OpenStudy (anonymous):

An angle measuring 3.5 radians is equal to which of the angle measures given below? Check all that apply. A. 210 degrees B. 200 degrees 32' 6'' C. 200.54 degrees D. 630 degrees

OpenStudy (kropot72):

Pi radians = 180 degrees The angle in decimal degrees is \[\frac{3.5}{\pi} \times 180\ degrees\]

OpenStudy (anonymous):

thank you :)

OpenStudy (kropot72):

You're welcome :)